OpenStudy (anonymous):

simplify sqrt 3x^8/sqrt 12x^2

OpenStudy (saifoo.khan):

\[\large \frac{\sqrt{3x+8}}{\sqrt{12x^2}}\]

OpenStudy (saifoo.khan):

like that?

OpenStudy (anonymous):

saif, shouldn't that "+" sign should be a "^" ?

OpenStudy (saifoo.khan):

LOL, right. TYPO.

OpenStudy (anonymous):

it is ^8 not +8

OpenStudy (saifoo.khan):

lol, i missed radicals! >.<

OpenStudy (anonymous):

@saif i thought you were good with them

OpenStudy (anonymous):

\[\frac{\sqrt{3x^8}}{\sqrt{12x^2}}=\frac{x^4\sqrt{3}}{x \sqrt{12}}=\frac{x^3}{\sqrt{4}}=\frac{x^3}{2}\]

OpenStudy (saifoo.khan):

\[\Huge \frac{x^3}{2} \]

OpenStudy (saifoo.khan):

This site is laging.

OpenStudy (saifoo.khan):

Ajmal's right.

OpenStudy (anonymous):

is that the correct answer

OpenStudy (saifoo.khan):

100%